Write the slope-intercept equation for the line that passes through (70, 92) and is (a) parallel and (b) perpendicular to 7x + 10y = -24. Please show all of your work.
Answer by mananth(16949)

PARALLEL


7 x + 10 y = -24
Find the slope of this line
make y the subject
10 y = -7 x + -24
Divide by 10
y = -0.7 x -2.4
Compare this equation with y=mx+b
slope m = -0.7 =>- 7/10
The slope of a line parallel to the above line will be the same m= -7/10 ,point ( 70 , 92 )
Find b by plugging the values of m & the point in
y=mx+b
92 = -49 + b
b= 141
m= -7/10
Plug value of the slope and b in y = mx +b
The required equation is y = -7/10 x + 141

---------------
PERPENDICULAR TO A LINE 7 x + 10 y = -24
Find the slope of this line

10 y = -7 x + 6
Divide by 10
y = -7/10 x + 3/5
Compare this equation with y=mx+b
slope m = -7/10

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 10/7

m= 10/7 ,point ( 70 , 92 )
Find b by plugging the values of m & the point in
y=mx+b
92 = 100.00 + b
b= -8
m= 10/7
The required equation is y = 10/ 7 x -8
m.ananth@hotmail.ca